#1 #3 } #6

#2

\overbrace{#5} #4

#8 #7

This is the type of technical illustration that draw2latex is good at:

draw2latex example file

[figure] [figure] [figure]

32

48

S2 S1 S3 S4 S5 S6 S7 S8

6 4

32

32 48

L R C

D

Expansion Permutation

[figure]

64 bit input 64 bit output

56 bit key

Initial permutation Final permutation Key permutation Key schedule text label

lines

beziérs

box grouped with text

[figure]

circle

[figure]

text area

Unfortunately, the TEX warnings with diagonal arrowheads cannot be avoided.

[figure]